E9-10 (Algo) Calculating Variable Overhead Variances [LO 9-5]

Parker Plastic, Inc., manufactures plastic mats to use with rolling office chairs. Its standard cost information for last year follows:

Standard Quantity Standard Price (Rate) Standard Unit Cost
Direct materials (plastic) 15 sq ft. $ 0.90 per sq. ft. $ 13.50
Direct labor 0.25 hr. $ 12.60 per hr. 3.15
Variable manufacturing overhead (based on direct labor hours) 0.25 hr. $ 1.80 per hr. 0.45
Fixed manufacturing overhead ($539,400 ÷ 930,000 units) 0.58

Parker Plastic had the following actual results for the past year:

Number of units produced and sold 1,240,000
Number of square feet of plastic used 12,400,000
Cost of plastic purchased and used $ 9,920,000
Number of labor hours worked 328,000
Direct labor cost $ 3,968,800
Variable overhead cost $ 1,350,000
Fixed overhead cost $ 385,000


Required:
Calculate Parker Plastic’s variable overhead rate and efficiency variances and its over- or underapplied variable overhead. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

Variable Overhead Rate Variance
Variable Overhead Efficiency Variance
Variable Overhead Spending Variance
Answer #1

Voh Rate variance = (SR-AR) *AH

= (1.80 - 1,350,000/328,000)*328,000

= 759,600 U

Voh efficiency variance = (SH-AH) *SR

= (1,240,000*0.25 - 328,000)*1.80

= 32,400 U

Voh spending variance = 759,600U + 32,400U

= 792,000U